# Fast ECoT: Efficient Embodied Chain-of-Thought via Thoughts Reuse

### v1

@@ 5.3 Real-world Experiments
Fast ECoT 76.6 66.7 46.7 63.3 2479 $\pm$ 520
Fast ECoT (Async) 83.3 76.7 50.0 70.0 716 $\pm$ 529
@@
Table 2: Real-world experimental results on selected household tasks. SR = Success Rate. ID = In distribution. OOD = Out of distribution. All results are averaged over 180 trials and run on a single NVIDIA RTX 4090 GPU. Best results are in bold.
Results. Tab. 2 summarises the performance of our method against baseline approaches on real-world tasks. Fast ECoT and Fast ECoT (Async) both demonstrate substantial latency reductions while preserving or even improving task success rates. Specifically, Fast ECoT (Async) delivers the best performance, with an average success rate of 70% across in-distribution (ID), out-of-distribution (OOD) objects, and OOD instruction tasks, alongside the lowest latency of $716\pm 529$ ms per step—a 7.5x speedup over the original ECoT baseline ($5556\pm 384$ ms). This result reinforces the benefits of asynchronous parallel reasoning in real-world deployment scenarios, where low-variance and rapid inference are critical.
5.4 Action Faithfulness in ECoT Reasoning
@@
Figure 6: Action faithfulness (AF) on LIBERO tasks. All graphs are plotted with mean and standard deviation shading across 1000 steps.
@@
To validate that the generated chain-of-thought (CoT) not only improves performance but also faithfully explains the model’s decision-making process, we build on the faithfulness criteria for language models [36] and introduce a novel quantitative metric - action faithfulness (AF) - to measure the faithfulness of the CoT reasoning for robotic tasks. fFormally, given a complete CoT consisting of $N$ intermediate reasoning steps that culminate in a final robot action $A$, we enforce the model to directly predict an action $A_{i}$ after generating only the first $i$ reasoning steps, where $i=0,1,\dots,N$. We then compute the L1 distance between $A_{i}$ and $A$ as the faithfulness score $AF_{i}=\|A_{i}-A\|_{1}$. Higher L1 distances indicate a greater dependence on subsequent reasoning steps, thereby suggesting higher CoT faithfulness.
@@
We plot action faithfulness scores of Fast ECoT and ECoT for all reasoning steps $\{AF_{i}\}$ in Fig. 6. Fast ECoT preserves the faithfulness of the base ECoT model during parallel reasoning generation for most reasoning steps. The action faithfulness without predicting “Visible Objects” is slightly lower for Fast ECoT, since the high update ratio and one-step delay in visual features might increase its chances to be post-hoc. Meanwhile, the asynchronous variant of Fast ECoT generally exhibits a lower faithfulness, probably due to the larger temporal mismatch introduced in async reasoning. Faithfulness is notably low when no reasoning steps are generated (i.e., at “Task”), likely due to fine-tuning from the OpenVLA [1] checkpoint, which did not require reasoning steps. We provide more CoT reasoning examples in Sec. C.1.
6 Conclusion
@@
7 Limitations
While our work demonstrated that ECoT reasoning can be significantly accelerated through the reuse of previously computed thoughts, there remain several limitations. First, we still update all reasoning steps at once without identifying and prioritising the more vital steps for frequent updates. This approach can lead to excessive computation, and more sophisticated methods could dynamically determine which reasoning steps require updating based on their relative importance to task success. Second, the effectiveness of thoughts reuse inherently relies on the assumption that at least part of the reasoning remains stable over a short time window. In scenarios where reasoning steps frequently change between consecutive time steps, this stability assumption breaks down, potentially degrading the performance gains offered by our method. Third, our current approach assumes a fixed structure in reasoning chains, thus limiting its applicability to tasks or environments requiring highly dynamic or adaptive reasoning patterns. Extending thoughts reuse to accommodate more diverse and adaptive reasoning structures remains an open challenge. Addressing these limitations represents promising avenues for future research to further enhance the efficiency and generality of embodied reasoning in robotic control.
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix A Extra Algorithm Details
A.1 Algorithm on Embodied Chain-of-Thought
Algorithm 2 Embodied Chain‑of‑Thought
1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$
2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
3:$R^{t}\leftarrow[\,]$
4:for $i=1\,\mathbf{to}\,N+1$ do
5:     Autoregressively sample $r_{i}^{t}\sim P\bigl{(}r_{i}^{t}\mid c^{t},R^{t}\bigr{)}$
6:     Add $r_{i}^{t}$ to $R^{t}$
7:end for
[…]
9:return $R^{t},A^{t}$
A.2 Algorithm on Parallel Embodied Chain-of-Thought
Algorithm 3 Parallel Embodied Chain‑of‑Thought
2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
3:$R^{t}\leftarrow[\,]$
4:for $i=1\,\mathbf{to}\,N+1$ concurrently do
6:end for
7:Synchronize
[…]
9:return $R^{t},A^{t}$
Appendix B Extra Experimental Setup
B.1 Finetuning Hyperparameters
Hyperparameter Value
Base Model ecot-openvla-7b-oxe
Number of GPUs 4
Batch Size 1
Gradient Accumulation Steps 1
Learning Rate 5e-4
Learning Rate Scheduler constant LR
Optimizer AdamW
AdamW Betas (0.9, 0.999)
AdamW Epsilon 1e-8
Training Steps 200,000
Image Augmentation
Reasoning Dropout 0.0
LoRA Rank 32
LoRA Alpha 16
LoRA Dropout 0.0
Table 3: Fine-tuning hyperparameters used for fine-tuning EcoT model.
B.2 Real-World Experimental Setup
Out-of-distribution Instructions. We use GPT-4o [37] to rephrase instructions from the training set. These rephrasings modify both the verbs and nouns to create linguistically novel yet semantically equivalent instructions, allowing us to test generalization to unseen language.
Original Instruction OOD Rephrased Instruction
Pick up the gray spanner and put it in the basket Grab the silver wrench and drop it into the bin
Pick up the blue hammer and put it in the basket Take the cobalt mallet and place it in the container
Pick up the blue squash ball and put it in the basket Retrieve the navy rubber ball and set it in the basket
Pick up the white baseball and put it in the basket Grab the pale softball and store it in the bucket
Pick up the banana and put it on the plate Take the yellow fruit and lay it on the dish
Pick up the lemon and put it in the cup Seize the citrus and drop it into the mug
Table 4: Original instructions and their corresponding out-of-distribution (OOD) rephrasings generated by GPT-4o.
In-distribution Objects.
Figure 7: Qualitative examples of in-distribution (ID) objects used for evaluation.
Out-of-distribution objects.
Figure 8: Qualitative examples of out-of-distribution (OOD) objects used for evaluation.
Appendix C Extra Experimental Results
C.1 CoT Reasoning Examples
Figure 9: Qualitative examples of ECOT reasoning in real-world experiments.
Figure 10: Qualitative examples of ECOT reasoning in simulation. As the VLMs used in both the data generation pipeline and the underlying VLA model are pretrained on real-world images, their performance in simulated environments is less stable. This can lead to inaccurate bounding boxes, undetected objects, and unusual or incorrect object labels.

### v2

@@ V-B Real-world Experiments
Fast ECoT 81.6 73.3 50.0 68.3 2479 $\pm$ 520
Fast ECoT (Async) 78.3 75.0 42.7 65.3 716 $\pm$ 529
@@
TABLE II: Real-world experimental results on selected household tasks. SR = Success Rate. ID = In distribution. OOD = Out of distribution.
Results. Tab. II summarises the performance of our method against baseline approaches on real-world manipulation tasks. Fast ECoT achieves the highest overall success rate (68.3%), with strong improvements in OOD instruction tasks (50.0%, +10 points over OpenVLA). Meanwhile, the Async version of Fast ECoT yields the best trade-off between performance and efficiency: although its average success rate is slightly lower (65.3%), it delivers the lowest latency of $716\pm 529$ ms per step—a $7.7\times$ speedup compared to the original ECoT baseline ($5556\pm 384$ ms). This highlights the benefit of asynchronous parallel reasoning in real-world deployment, where rapid, low-variance inference is critical. In contrast, prior ECoT variants either sacrifice task performance (ECoT 5-step) or still suffer from high inference latency (ECoT and ECoT-Async), making them less practical for on-robot use. Consistent with the simulation examples, Fig. 8 shows that both methods produce coherent, grounded reasoning; again, the asynchronous variant reaches comparable task states much quicker.
(a) ECoT at $t=84.8$s.
(b) Fast ECoT (Async) at $t=18.7$s.
Figure 8: Real-world qualitative examples. Fast ECoT (Async) reaches comparable task states much quicker.
V-C AutoEval Real-world Experiments
Setup. We use AutoEval [45], an online platform for standardised real-robot policy evaluation. We choose AutoEval both for reproducible, third-party benchmarking and because its tasks/environments are drawn from BridgeData V2 [5]—the same distribution on which ECoT [4] was trained. This lets us run the vanilla ECoT without any finetuning and directly apply our accelerations (Fast ECoT and Fast ECoT (Async)) as drop-in replacements. Concretely, we evaluate four tabletop tasks in two accessible environment (Drawer, Sink) with 10 trials per task.
Method Drawer SR (%) $\uparrow$ Sink SR (%) $\uparrow$ Latency (ms) $\downarrow$
ECoT 30 30 4030 $\pm$ 270
Fast ECoT 30 35 2105 $\pm$ 324
Fast ECoT (Async) 20 25 790 $\pm$ 331
TABLE III: AutoEval results on BridgeData V2 tabletop tasks. SR = Success Rate.
Results. Fast ECoT and Fast ECoT (Async) achieve performance comparable to ECoT under AutoEval (Tab. III): Fast ECoT matches ECoT on Drawer (30%) and slightly exceeds it on Sink (35%), while the asynchronous variant is modestly lower (20–25%) but offers a favourable accuracy–efficiency trade-off. Note that AutoEval’s environment reset policy introduces distribution shift—e.g., altered object poses/orientations, handle angles, and initial camera/scene configurations—relative to the canonical starts used in the original evaluation. These OOD initialisations make perception and short-horizon control more volatile, reducing absolute SRs across all methods (including ECoT) even though their relative behaviour remains similar.
V-D Updating Frequency of Across-level Reasoning
Reusing past reasoning steps yields a temporal smoothing effect that dampens noisy fluctuations in planning and stabilises control. We perform a controlled ablation on LIBERO-Object (see Tab. IV) to examine performance differences resulting from varying the update frequency of high-level and low-level reasoning steps. When high-level reasoning is refreshed less frequently (every 5 frames), while low-level reasoning remains updated at every step, success rates increase from $77\%$ (baseline with updates at every step) to $79\%$. In Fast ECoT, each reasoning step similarly leverages current visual observations combined with cached reasoning from previous steps, achieving an even higher success rate of $83\%$. However, excessively infrequent updates reduce performance. Updating both high-level and low-level reasoning every 5 frames lowers success to $70\%$, and completely infrequent updates ($\infty$$\infty$) dramatically drop success to $35\%$. Common failures in such scenarios involve delayed task transitions—for instance, after grasping an object, the robot may stall indefinitely rather than moving promptly to the placement step. This underscores the critical balance between beneficial temporal smoothing and necessary action responsiveness.
High-level step Low-level step Success (%) $\uparrow$
1 1 77
5 1 79
5 2 78
5 3 75
5 5 70
$\infty$ $\infty$ 35
Fast ECoT Fast ECoT 83
TABLE IV: Ablation on update frequency.
Our findings (see Tab. IV) align with the results in SmolVLA [37] and Tab. 2 of the original ECoT work, where the reuse of historical reasoning similarly improved task performance. Thus, moderate reuse of recent reasoning can maintain or even enhance robot performance.
V-E Action Faithfulness in ECoT Reasoning
@@
Figure 9: Action faithfulness (AF) on LIBERO tasks. All graphs are plotted with mean and standard deviation shading across 1000 timesteps.
@@
To validate that the generated chain-of-thought (CoT) not only improves performance but also faithfully explains the model’s decision-making process, we build on the faithfulness criteria for language models [46] and introduce a novel quantitative metric - action faithfulness (AF) - to measure the faithfulness of the CoT reasoning for robotic tasks. Formally, given a complete CoT consisting of $N$ intermediate reasoning steps that culminate in a final robot action $A$, we enforce the model to directly predict an action $A_{i}$ after generating only the first $i$ reasoning steps, where $i=0,1,\dots,N$. We then compute the L1 distance between $A_{i}$ and $A$ as the faithfulness score $AF_{i}=\|A_{i}-A\|_{1}$. Higher L1 distances indicate a greater dependence on subsequent reasoning steps, thereby suggesting higher CoT faithfulness.
@@
We plot action faithfulness scores of Fast ECoT and ECoT for all reasoning steps $\{AF_{i}\}$ in Fig. 9. Fast ECoT preserves the faithfulness of the base ECoT model during parallel reasoning generation for most reasoning steps. The action faithfulness without predicting “Visible Objects” is slightly lower for Fast ECoT, since the high update ratio and one-step delay in visual features might increase its chances to be post-hoc. Meanwhile, the asynchronous variant of Fast ECoT generally exhibits a lower faithfulness, probably due to the larger temporal mismatch introduced in async reasoning. Faithfulness is notably low when no reasoning steps are generated (i.e., at “Task”), likely due to fine-tuning from the OpenVLA [1] checkpoint, which did not require reasoning steps.
VI CONCLUSION